# Totally geodesic submanifolds of the homogeneous nearly Kähler 6-manifolds and their $\mathsf{G}_{2}$-cones

### v2

@@ 3.1. Maximal totally geodesic submanifolds in analytic Riemannian manifolds
We now concern ourselves with defining a notion of maximality for totally geodesic submanifolds. Indeed, if $M$ is a real analytic Riemannian manifold and $\Sigma_{1}$, $\Sigma_{2}\subseteq M$ are two inextendable and embedded totally geodesic submanifolds, one can ask if $\Sigma_{1}\subseteq\Sigma_{2}$. In this case, it is easy to see that $\Sigma_{1}\subseteq\Sigma_{2}$ if and only if there exists a point $p\in\Sigma_{1}\cap\Sigma_{2}$ such that $T_{p}\Sigma_{1}\subseteq T_{p}\Sigma_{2}$. Therefore, the study of inclusions between embedded totally geodesic submanifolds of $M$ containing the point $p$ is equivalent to that of inclusions between totally geodesic subspaces of $T_{p}M$. For general totally geodesic immersions, the situation is more involved, and one needs to introduce the following “pullback-type” construction to make sense of the inclusion relationship.
@@
$x\mathcal{R}y\colon\Leftrightarrow\widetilde{f_{2}\circ h}(x)=\widetilde{f_{2}\circ h}(y).$
@@ 4. Totally geodesic submanifolds in naturally reductive homogeneous spaces
In this section, we introduce new techniques for studying totally geodesic submanifolds in naturally reductive homogeneous spaces. Because homogeneous spaces are complete and real analytic, by Corollary 3.5 their inextendable totally geodesic submanifolds are complete. These submanifolds are also homogeneous as Riemannian manifolds as a consequence of [39, Theorem 8.9], but they need not be extrinsically homogeneous.
@@
We now give a geometric interpretation of the subspace $e^{-tD_{X}}\mathfrak{v}$. Consider the geodesic $\gamma(t)=\operatorname{Exp}(tX)\cdot o$ with initial condition $X\in\mathfrak{v}$. Then there are two vector space isomorphisms that we can establish between $\mathfrak{p}=T_{o}M$ and $T_{\gamma(t)}M$: parallel translation $\mathcal{P}_{0,t}^{\gamma}\colon T_{o}M\to T_{\gamma(t)}M$ and the pushforward of the flow of $X^{*}$, given by $\operatorname{Exp}(tX)_{*o}\colon T_{o}M\to T_{\gamma(t)}M$. Both maps are related by [46, Equation (2.2.1)]
@@
Suppose that $\mathfrak{v}$ is totally geodesic. We consider the complete totally geodesic immersion $f\colon\Sigma\to M$ such that $\mathfrak{v}=\widetilde{f}(p)$ and take $v=(f_{*p})^{-1}(X)\in T_{p}\Sigma$, $g=\operatorname{Exp}(-tX)$. From (5) and the fact that $f$ commutes with parallel translations we see that $e^{-tD_{X}}\mathfrak{v}=\widetilde{g\circ f}(\exp_{p}(tv))$, yielding the following result:
Corollary 4.2 ([55, Proposition 3.5]).
@@ 4.1. Totally geodesic submanifolds invariant under $D$
The following theorem gives an algebraic characterization of $D$-invariant totally geodesic submanifolds passing through the origin. Furthermore, it gives an explicit method to construct them from their tangent space at $o$. The proof can be obtained by combining the theorem in [51, p. 11] and the first result in [30, §2]. However, we will include it for the sake of completeness.
@@
Furthermore, if any of the three previous conditions hold, we have:
@@
Therefore, $\mathfrak{s}$ is a Lie subalgebra. Because $\mathfrak{v}$ is $D$-invariant, se wee that $\mathfrak{s}_{\mathfrak{p}}=\mathfrak{v}\subseteq\mathfrak{s}$, and from this inclusion it follows that $\mathfrak{s}=\mathfrak{s}_{\mathfrak{k}}\oplus\mathfrak{v}$, which proves that $\mathfrak{s}$ is canonically embedded. It is also immediate from the description of $\mathfrak{s}$ that $\mathfrak{s}_{\mathfrak{k}}=[\mathfrak{v},\mathfrak{v}]_{\mathfrak{k}}$. As a consequence, if we consider the Lie subgroup $\mathsf{S}$ of $\mathsf{G}$ whose Lie algebra is $\mathfrak{s}$, then the tangent space $T_{o}(\mathsf{S}\cdot o)$ coincides with $\mathfrak{s}_{\mathfrak{p}}=\mathfrak{v}$, and therefore (iii) holds.
We now prove that (iii) implies (iv). Assume $\mathsf{S}\subseteq\mathsf{G}$ is a Lie subgroup whose Lie algebra is canonically embedded in $\mathfrak{g}$ and $T_{o}(\mathsf{S}\cdot o)=\mathfrak{s}_{\mathfrak{p}}=\mathfrak{v}$. It is clear from (2) that $\mathbb{II}$ is zero at $o$. Because $\mathsf{S}\cdot o$ is extrinsically homogeneous, this implies that the second fundamental form vanishes everywhere, and thus $\mathsf{S}\cdot o$ is totally geodesic. We now prove that $\mathsf{S}\cdot o$ is $D$-invariant, which by $\mathsf{G}$-invariance of $D$ is equivalent to checking that $\mathfrak{v}$ is $D$-invariant. Given $X$, $Y\in\mathfrak{s}_{\mathfrak{p}}\subseteq\mathfrak{s}$, we have $D_{X}Y=(1/2)[X,Y]_{\mathfrak{p}}\in\mathfrak{s}_{\mathfrak{p}}$, so the claim follows.
@@
Every $D$-invariant totally geodesic submanifold of a naturally reductive homogeneous space $M=\mathsf{G}/\mathsf{K}$ is extrinsically homogeneous with respect to the given presentation of $M$, and thus an injectively immersed submanifold.
@@
Theorem 4.3 is also a refinement of [48, Lemma 3.1], which states that for a general reductive homogeneous space a subspace $\mathfrak{v}\subseteq\mathfrak{p}$ invariant under $R$ and $D$ is tangent to a complete totally geodesic submanifold. Also, notice that the class of $D$-invariant totally geodesic submanifolds includes all totally geodesic submanifolds of symmetric spaces, since in a symmetric space $D=0$, thus, in the symmetric setting the subspaces $\mathfrak{v}$ appearing in the previous theorem are the Lie triple systems (see [6, §11.1]).
@@
It is worth noting that in the irreducible setting the conditions of Theorem 4.3 do not depend on the naturally reductive decomposition that we choose. Indeed, from [47, Theorem 2.1], we see that if $M=\mathsf{G}/\mathsf{K}$ is a simply connected irreducible naturally reductive space which is not symmetric, then the canonical connection $\nabla^{c}$ is unique. Therefore, given any naturally reductive decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ of $\mathfrak{g}$, the subspaces $\mathfrak{v}\subseteq\mathfrak{p}$ that are invariant under $R$ and $D$ correspond under the identification $\mathfrak{p}\equiv T_{o}M$ to the subspaces $V\subseteq T_{o}M$ that are invariant under $R$ and $\nabla-\nabla^{c}$, and the uniqueness of the canonical connection implies that these subspaces are always the same regardless of the decomposition. Similarly, if $M$ is a nearly Kähler 3-symmetric space and one restricts their attention to the reductive decompositions invariant under the automorphism of order three, then all of their associated canonical connections coincide by [56, Lemma 3.1], and the same argument applies.
@@ 4.2. Totally geodesic surfaces
Let $\mathfrak{v}\subseteq\mathfrak{p}$ be a $2$-dimensional subspace, and assume that $\mathfrak{v}$ is the tangent plane at $o$ of the complete totally geodesic surface $\Sigma$ of $M$. Fix a nonzero element $X\in\mathfrak{v}$ and choose any $Y\in\mathfrak{v}\setminus\{0\}$ that is orthogonal to $X$, so that $\{X,Y\}$ is an orthogonal basis of $\mathfrak{v}$. Since $\Sigma$ is homogeneous and two-dimensional, it follows that $\Sigma$ is a space of constant curvature $\kappa\in\mathbb{R}$, and the same can be said for the totally geodesic submanifold $\Sigma_{t}$ associated with $e^{-tD_{X}}\mathfrak{v}$ for all $t\in\mathbb{R}$. This implies in particular that the restriction of $\nabla R$ to the tangent space of $\Sigma$ and $\Sigma_{t}$ at any of their points is the zero tensor. Furthermore, we have $e^{-tD_{X}}\mathfrak{v}=\operatorname{span}\{X,e^{-tD_{X}}Y\}$, since $D_{X}X=0$ due to the skew-symmetry of $D$. Because $\Sigma_{t}$ has curvature $\kappa$, it follows that $e^{-tD_{X}}Y$ is an eigenvector of the Jacobi operator $R_{X}$ with eigenvalue $\kappa\lvert X\rvert^{2}$, as well as an element of $\ker C_{X}$. One can argue similarly with the so-called Cartan operators of order $j$ given by $C^{j}_{X}Y=\nabla^{j}R(X,\dots,X,Y,X)$, because they vanish identically on $\mathfrak{v}$. Since the subspace of $\mathfrak{p}$ generated by the curve $e^{-tD_{X}}Y$ is the span of all vectors of the form $D_{X}^{k}Y$ with $k\geq 0$, we have obtained the following:
@@ 4.3. Well-positioned totally geodesic submanifolds and homogeneous fibrations
We now study the case that $M=\mathsf{G}/\mathsf{H}$ is also the total space of the homogeneous fibration induced by the inclusions $\mathsf{H}\subseteq\mathsf{K}\subseteq\mathsf{G}$ (observe the change of notation). Let $B=\mathsf{G}/\mathsf{K}$ be the base space and $F=\mathsf{K}/\mathsf{H}$ be the fiber of the given submersion. Consider a totally geodesic immersion $f\colon\Sigma\to M$. We say that $\Sigma$ is well-positioned at $p\in\Sigma$ (with respect to the fibration $M\to B$) if $\widetilde{f}(p)=\left(\widetilde{f}(p)\cap\mathcal{V}_{f(p)}\right)\oplus\left(\widetilde{f}(p)\cap\mathcal{H}_{f(p)}\right)$. Furthermore, $\Sigma$ is said to be well-positioned if it is well-positioned at every point $p\in\Sigma$. The next result allows us to give an algebraic characterization for a totally geodesic submanifold to be well-positioned.
@@
Let $F\to M\to B$ be the homogeneous fibration induced by the chain of inclusions $\mathsf{H}\subseteq\mathsf{K}\subseteq\mathsf{G}$, where $\mathsf{H}$, $\mathsf{K}$ and $\mathsf{G}$ are compact and the Riemannian metrics on $F$, $M$ and $B$ are induced from a bi-invariant metric on $\mathsf{G}$. Let $f\colon\Sigma\to M$ be a complete totally geodesic immersion passing through the point $o$ with tangent space $\mathfrak{v}$. The following conditions are equivalent:
@@
Let $p\in\Sigma$ be such that $\widetilde{f}(p)=\mathfrak{v}$. Since we are assuming that $\Sigma$ is connected and complete, every point of $\Sigma$ is of the form $q=\exp_{p}(v)$ for a certain $v\in T_{p}\Sigma$. Consider the geodesic $\gamma(t)=f(\exp(tv))=\operatorname{Exp}(tX)\cdot o$ (where $X=f_{*p}(v)$) connecting $p$ and $q$. Then, we have $\widetilde{f}(q)=\mathcal{P}_{0,1}^{\gamma}\mathfrak{v}$. As $\mathcal{V}$ and $\mathcal{H}$ are invariant under $\mathsf{G}$, we see that $\mathcal{V}_{f(q)}=\operatorname{Exp}(X)_{*o}\mathcal{V}_{o}$ and $\mathcal{H}_{f(q)}=\operatorname{Exp}(X)_{*o}\mathcal{H}_{o}$. Therefore, using (5) we see that $\Sigma$ is well-positioned at $q$ if and only if the subspace $\operatorname{Exp}(X)_{*o}^{-1}(\widetilde{f}(q))=e^{-D_{X}}\mathfrak{v}$ splits with respect to the decomposition $\mathfrak{p}=\mathcal{V}_{o}\oplus\mathcal{H}_{o}$. As $\mathfrak{v}=f_{*p}(T_{p}\Sigma)$, the equivalence follows. ∎

### v3

@@ 3.1. Maximal totally geodesic submanifolds in analytic Riemannian manifolds
We now concern ourselves with defining a notion of maximality for totally geodesic submanifolds. Indeed, if $M$ is a real analytic Riemannian manifold and $\Sigma_{1}$, $\Sigma_{2}\subseteq M$ are two inextendable and embedded totally geodesic submanifolds, one can wonder if $\Sigma_{1}\subseteq\Sigma_{2}$. In this case, it is easy to see that $\Sigma_{1}\subseteq\Sigma_{2}$ if and only if there exists a point $p\in\Sigma_{1}\cap\Sigma_{2}$ such that $T_{p}\Sigma_{1}\subseteq T_{p}\Sigma_{2}$. Therefore, the study of inclusions between embedded totally geodesic submanifolds of $M$ containing the point $p$ is equivalent to that of inclusions between totally geodesic subspaces of $T_{p}M$. For general totally geodesic immersions, the situation is more involved, and one needs to introduce the following “pullback-type” construction to make sense of the inclusion relationship.
@@
$x\mathcal{R}y:\Leftrightarrow\widetilde{f_{2}\circ h}(x)=\widetilde{f_{2}\circ h}(y).$
@@ 4. Totally geodesic submanifolds in naturally reductive homogeneous spaces
In this section, we introduce new techniques for studying totally geodesic submanifolds in naturally reductive homogeneous spaces. Because homogeneous spaces are complete and real analytic, by Corollary 3.5 their inextendable totally geodesic submanifolds are complete. These submanifolds are also homogeneous as Riemannian manifolds as a consequence of [38, Theorem 8.9], but they need not be extrinsically homogeneous.
@@
We now give a geometric interpretation of the subspace $e^{-tD_{X}}\mathfrak{v}$. Consider the geodesic $\gamma(t)=\operatorname{Exp}(tX)\cdot o$ with initial condition $X\in\mathfrak{v}$. Then there are two vector space isomorphisms that we can establish between $\mathfrak{p}=T_{o}M$ and $T_{\gamma(t)}M$: parallel translation $\mathcal{P}_{0,t}^{\gamma}\colon T_{o}M\to T_{\gamma(t)}M$ and the pushforward of the flow of $X^{*}$, given by $\operatorname{Exp}(tX)_{*o}\colon T_{o}M\to T_{\gamma(t)}M$. Both maps are related by [50, Equation (2.2.1)]
@@
Suppose that $\mathfrak{v}$ is totally geodesic. We consider the complete totally geodesic immersion $f\colon\Sigma\to M$ such that $\mathfrak{v}=\widetilde{f}(p)$ for some point $p\in\Sigma$ and take $v=(f_{*p})^{-1}(X)\in T_{p}\Sigma$, $g=\operatorname{Exp}(-tX)$. From (5) and the fact that $f$ commutes with parallel translations we see that $e^{-tD_{X}}\mathfrak{v}=\widetilde{g\circ f}(\exp_{p}(tv))$, yielding the following result:
Corollary 4.2 ([54, Proposition 3.5]).
@@ 4.1. Totally geodesic submanifolds invariant under $D$
The following theorem gives an algebraic characterization of $D$-invariant totally geodesic submanifolds passing through the origin. Furthermore, it gives an explicit method to construct them from their tangent space at $o$. The proof can be obtained by combining the theorem in [49, p. 11] and the first result in [29, §2]. However, we include it for the sake of completeness.
@@
Furthermore, if any of the four previous conditions hold, we have:
@@
Therefore, $\mathfrak{s}$ is a Lie subalgebra. Because $\mathfrak{v}$ is $D$-invariant, we see that $\mathfrak{s}_{\mathfrak{p}}=\mathfrak{v}\subseteq\mathfrak{s}$, and from this inclusion it follows that $\mathfrak{s}=\mathfrak{s}_{\mathfrak{k}}\oplus\mathfrak{v}$, which proves that $\mathfrak{s}$ is canonically embedded. It is also immediate from the description of $\mathfrak{s}$ that $\mathfrak{s}_{\mathfrak{k}}=[\mathfrak{v},\mathfrak{v}]_{\mathfrak{k}}$. As a consequence, if we consider the Lie subgroup $\mathsf{S}$ of $\mathsf{G}$ whose Lie algebra is $\mathfrak{s}$, then the tangent space $T_{o}(\mathsf{S}\cdot o)$ coincides with $\mathfrak{s}_{\mathfrak{p}}=\mathfrak{v}$, and therefore (iii) holds.
We now prove that (iii) implies (iv). Assume $\mathsf{S}\subseteq\mathsf{G}$ is a Lie subgroup whose Lie algebra is canonically embedded in $\mathfrak{g}$ and $T_{o}(\mathsf{S}\cdot o)=\mathfrak{s}_{\mathfrak{p}}=\mathfrak{v}$. It is clear from (2) that $\mathbb{II}$ is zero at $o$. Because $\mathsf{S}\cdot o$ is extrinsically homogeneous, this implies that the second fundamental form vanishes everywhere, and thus $\mathsf{S}\cdot o$ is totally geodesic. We show that $\mathsf{S}\cdot o$ is $D$-invariant, which by $\mathsf{G}$-invariance of $D$ is equivalent to checking that $\mathfrak{v}$ is $D$-invariant. Given $X$, $Y\in\mathfrak{s}_{\mathfrak{p}}\subseteq\mathfrak{s}$, we have $D_{X}Y=(1/2)[X,Y]_{\mathfrak{p}}\in\mathfrak{s}_{\mathfrak{p}}$, so the claim follows.
@@
Every complete $D$-invariant totally geodesic submanifold of a naturally reductive homogeneous space $M=\mathsf{G}/\mathsf{K}$ is extrinsically homogeneous with respect to the given presentation of $M$, and thus an injectively immersed submanifold.
@@
Theorem 4.3 is also a refinement of [46, Lemma 3.1], which states that for a general reductive homogeneous space a subspace $\mathfrak{v}\subseteq\mathfrak{p}$ invariant under $R$ and $D$ is tangent to a complete totally geodesic submanifold. Also, notice that the class of $D$-invariant totally geodesic submanifolds includes all totally geodesic submanifolds of symmetric spaces, since in a symmetric space $D=0$. Thus, in the symmetric setting the subspaces $\mathfrak{v}$ appearing in the previous theorem are the Lie triple systems (see [6, §11.1]).
@@
It is worth noting that in the irreducible setting the conditions of Theorem 4.3 do not depend on the naturally reductive decomposition that we choose. Indeed, from [45, Theorem 2.1], we see that if $M=\mathsf{G}/\mathsf{K}$ is a simply connected irreducible naturally reductive space which is not symmetric, then the canonical connection $\nabla^{c}$ is unique. Therefore, given any naturally reductive decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ of $\mathfrak{g}$, the subspaces $\mathfrak{v}\subseteq\mathfrak{p}$ that are invariant under $R$ and $D$ correspond under the identification $\mathfrak{p}\equiv T_{o}M$ to the subspaces $V\subseteq T_{o}M$ that are invariant under $R$ and $\nabla-\nabla^{c}$, and the uniqueness of the canonical connection implies that these subspaces are always the same regardless of the decomposition. Similarly, if $M$ is a nearly Kähler 3-symmetric space and one restricts their attention to the reductive decompositions invariant under the automorphism of order three, then all of their associated canonical connections coincide by [55, Lemma 3.1], and the same argument applies.
@@ 4.2. Totally geodesic surfaces
Let $\mathfrak{v}\subseteq\mathfrak{p}$ be a $2$-dimensional subspace, and assume that $\mathfrak{v}$ is the tangent plane at $o$ of a complete totally geodesic surface $\Sigma$ of $M$. Fix a nonzero element $X\in\mathfrak{v}$ and choose any $Y\in\mathfrak{v}\setminus\{0\}$ that is orthogonal to $X$, so that $\{X,Y\}$ is an orthogonal basis of $\mathfrak{v}$. Since $\Sigma$ is homogeneous and two-dimensional, it follows that $\Sigma$ is a space of constant curvature $\kappa\in\mathbb{R}$, and the same can be said for the totally geodesic submanifold $\Sigma_{t}$ associated with $e^{-tD_{X}}\mathfrak{v}$ for all $t\in\mathbb{R}$. This implies in particular that the restriction of $\nabla R$ to the tangent space of $\Sigma$ and $\Sigma_{t}$ at any of their points is the zero tensor. Furthermore, since $D_{X}X=0$ due to the skew-symmetry of $D$, we have $e^{-tD_{X}}\mathfrak{v}=\operatorname{span}\{X,e^{-tD_{X}}Y\}$. Because $\Sigma_{t}$ has curvature $\kappa$, it follows that $e^{-tD_{X}}Y$ is an eigenvector of the Jacobi operator $R_{X}$ with eigenvalue $\kappa\lvert X\rvert^{2}$, as well as an element of $\ker C_{X}$. One can argue similarly with the so-called Cartan operators of order $j$ given by $C^{j}_{X}Y=\nabla^{j}R(X,\dots,X,Y,X)$, because they vanish identically on $\mathfrak{v}$. Since the subspace of $\mathfrak{p}$ generated by the curve $e^{-tD_{X}}Y$ is the span of all vectors of the form $D_{X}^{k}Y$ with $k\geq 0$, we have obtained the following:
@@ 4.3. Well-positioned totally geodesic submanifolds and homogeneous fibrations
We now study the case that $M=\mathsf{G}/\mathsf{H}$ is also the total space of the homogeneous fibration induced by the inclusions $\mathsf{H}\subseteq\mathsf{K}\subseteq\mathsf{G}$ (observe the change of notation). Let $B=\mathsf{G}/\mathsf{K}$ be the base space and $F=\mathsf{K}/\mathsf{H}$ be the fiber of the given submersion. Consider a totally geodesic immersion $f\colon\Sigma\to M$. We say that $\Sigma$ is well-positioned at $p\in\Sigma$ (with respect to the fibration $F\to M\to B$) if $\widetilde{f}(p)=\left(\widetilde{f}(p)\cap\mathcal{V}_{f(p)}\right)\oplus\left(\widetilde{f}(p)\cap\mathcal{H}_{f(p)}\right)$. Furthermore, $\Sigma$ is said to be well-positioned if it is well-positioned at every point $p\in\Sigma$. The next result allows us to give an algebraic characterization for a totally geodesic submanifold to be well-positioned.
@@
Let $F\to M\to B$ be the homogeneous fibration induced by the chain of inclusions $\mathsf{H}\subseteq\mathsf{K}\subseteq\mathsf{G}$, where $\mathsf{H}$, $\mathsf{K}$ and $\mathsf{G}$ are compact and the Riemannian metrics on $F$, $M$ and $B$ are induced by a bi-invariant metric on $\mathsf{G}$. Let $f\colon\Sigma\to M$ be a complete totally geodesic immersion passing through the point $o$ with tangent space $\mathfrak{v}$. Then, the following conditions are equivalent:
@@
Let $p\in\Sigma$ be such that $\widetilde{f}(p)=\mathfrak{v}$. Since we are assuming that $\Sigma$ is connected and complete, every point of $\Sigma$ is of the form $q=\exp_{p}(v)$ for a certain $v\in T_{p}\Sigma$. Consider the geodesic $\gamma(t)=f(\exp(tv))=\operatorname{Exp}(tX)\cdot o$ (where $X=f_{*p}(v)$), which connects $p$ and $q$. Then, we have $\widetilde{f}(q)=\mathcal{P}_{0,1}^{\gamma}\mathfrak{v}$. As $\mathcal{V}$ and $\mathcal{H}$ are invariant under $\mathsf{G}$, we see that $\mathcal{V}_{f(q)}=\operatorname{Exp}(X)_{*o}\mathcal{V}_{o}$ and $\mathcal{H}_{f(q)}=\operatorname{Exp}(X)_{*o}\mathcal{H}_{o}$. Therefore, using (5) we see that $\Sigma$ is well-positioned at $q$ if and only if the subspace $\operatorname{Exp}(X)_{*o}^{-1}(\widetilde{f}(q))=e^{-D_{X}}\mathfrak{v}$ splits with respect to the decomposition $\mathfrak{p}=\mathcal{V}_{o}\oplus\mathcal{H}_{o}$. As $\mathfrak{v}=f_{*p}(T_{p}\Sigma)$, the equivalence follows. ∎